Circular No. 8810

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

ANOTHER POSSIBLE NOVA IN SCORPIUS H. Yamaoka, Kyushu University, reports the discovery by Yuji Naka-

mura (Kameyama, Mie, Japan) of a variable star (mag 9.3) on two unfiltered CCD images taken with a 135-mm lens on Feb. 19.8593 UT (not present to mag 12.0 on Feb. 14.8575); the position was stated as being within 30′′ of the red star GSC 7367.712 (α = 16^{h}56^{m}58^{s}.0, δ = −35^{o}21′39′′, equinox 2000.0, mag 13.3; variable according to ASAS). K. Itagaki (Yamagata, Japan) obtained an unfiltered CCD image with a 0.60-m reflector on Feb. 20.85, concluding that a new star of mag 8.8 is located at α = 16^{h}56^{m}59^{s}.35, δ = −35^{o}21′50′′.2; Itagaki notes that nothing is visible at this position on the Digitized Sky Survey, so it is likely to be a nova. S. Nakano, Sumoto, Japan, reports the independent discovery of this variable by Hideo Nishimura (Miyawaki, Kakegawa) at mag 9.2 on two T-Max 400 films taken on Feb. 20.8365 and 20.8370 using a Pentax 6×7 camera (the position end figures given as 55^{s}, 59′′), with nothing visible at this location on his survey film taken on Feb. 15.8 (limiting mag 11.0).

COMET C/2007 D3 (LINEAR) An apparently asteroidal object discovered by the LINEAR project

(discovery observation tabulated below) has been found to show cometary appearance following posting on the Minor Planet Center’s ‘NEOCP’ webpage. Thirty stacked 75-s CCD exposures taken on Feb. 20.94–20.96 UT by C. Jacques and E. Pimentel (Belo Horizonte, Minas Gerais, Brazil, 0.30-m f/3 Schmidt-Cassegrain telescope) show a central condensation 5′′ in diameter and a small fan tail 15′′ long in p.a. 98^{o}. J. E. McGaha (Tucson, AZ, U.S.A., 0.36-m f/10 Schmidt-Cassegrain reflector) writes that his CCD images from Feb. 21.2 show a 4′′ coma surrounding a faint, stellar, nuclear condensation, with no elongation or tail visible. R. E. Hill (Mount Lemmon 1.5-m reflector) reports that his stack of four unfiltered 30-s CCD frames taken on Feb. 21.2 (in 2′′–2′′.5 seeing) show a coma diameter of 10′′–12′′ and again no tail.

2007 UT  α_{2000}  δ_{2000}  Mag.
Feb. 20.09455  5^{h}36^{m}56.^{s}96  −19^{o}37′15′′.5  19.2

The available astrometry, preliminary parabolic orbital elements (T = 2007 May 26.804 TT, q = 5.22363 AU, ω = 309^{o}.100, Ω = 148^{o}.199, i = 46^{o}.004, equinox 2000.0), and an ephemeris appear on MPEC 2007-D30.

2007 February 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green